# Relaxed Contrastive Learning for Federated Learning

### v1

@@ Abstract
We propose a novel contrastive learning framework to effectively address the challenges of data heterogeneity in federated learning. We first analyze the inconsistency of gradient updates across clients during local training and establish its dependence on the distribution of feature representations, leading to the derivation of the supervised contrastive learning (SCL) objective to mitigate local deviations. In addition, we show that a naïve adoption of SCL in federated learning leads to representation collapse, resulting in slow convergence and limited performance gains. To address this issue, we introduce a relaxed contrastive learning loss that imposes a divergence penalty on excessively similar sample pairs within each class. This strategy prevents collapsed representations and enhances feature transferability, facilitating collaborative training and leading to significant performance improvements. Our framework outperforms all existing federated learning approaches by huge margins on the standard benchmarks through extensive experimental results. We plan to release the source code of our work for better reproducibility.
@@ 1 Introduction
Federated learning (FL) trains a shared model through the collaboration of distributed clients while safeguarding the privacy of local data by restricting their sharing and transfer. The primary challenge in this learning framework arises from the data heterogeneity across clients and the class imbalance in local data. These problems eventually lead to severe misalignments of the local optima of the client models, hindering the search for better global optima of the aggregated model and slowing down the convergence.
In addressing these challenges, most existing approaches focus on minimizing the discrepancy between the global and local models by incorporating regularization techniques on either model parameters [21, 12, 2, 44, 1] or feature representations [19, 23, 18, 40, 15], aiming to balance between local training and regularization. However, aligning the local models with the global model entails a trade-off as the global model is not necessarily optimal. Recently, there have been several attempts to analyze the inconsistent local training in a principled way [42, 33]. For example, Zhang et al. [42] investigate the label distribution skewness from a statistical perspective, and introduce a deviation bound for analyzing the inconsistency of gradient updates in local training.
We reformulate the deviation bound proposed in [42] and establish its dependence on the distribution of feature representations. Subsequently, we derive that incorporating the supervised contrastive learning (SCL) objective enhances this bound, resulting in consistent local updates across heterogeneous clients. In other words, we show that employing SCL facilitates federated learning by alleviating the variations of local models.
(a) Tiny-ImageNet
@@
Figure 1: Performance curves of our framework, dubbed as FedRCL, in comparison to other baselines on the Tiny-ImageNet and CIFAR-100 with non-i.i.d. setting ($\alpha=0.1$). FedSCL incorporates the supervised contrastive learning objective into FedAvg, but it suffers from slow convergence and restrained performance enhancement. Our framework is significantly superior in both convergence speed and performance improvement.
Although SCL is helpful for the optimization in federated learning, the empirical results show that a naïve integration of SCL suffers from slow convergence and limited performance gains as shown in Figure 1. Due to the limited and imbalanced data in the local training set, the intra-class attraction force in SCL hampers feature diversity and consequently weakens the transferability of neural networks to diverse tasks. Considering FL’s nature of ongoing updates through the aggregation of heterogeneous local models, the lack of transferability impedes the collaborative training process.
To tackle this feature collapse issue and enhance the transferability of models, we present a novel contrastive learning technique for federated learning. Our approach imposes the penalty on the sample pairs within the same class that exhibit excessively high similarity. Such a simple adaptive repulsion strategy effectively prevents the intra- and inter-class collapse of representations, enhancing the transferability across heterogeneous clients and leading to the discovery of better global optima. Furthermore, we expand the proposed approach to cover all intermediate levels of representations, promoting consistent local updates even further. In spite of its simplicity, the proposed approach demonstrates remarkable performance improvements in all datasets and settings consistently, surpassing existing baselines by substantial margins. We show the effectiveness and robustness of the proposed method by thorough empirical analysis. Our main contributions are summarized as follows.
  • $\bullet$
    Based on our deviation bound, we theoretically analyze that supervised contrastive learning mitigates inconsistent local updates across heterogeneous clients.
  • $\bullet$
    We discover the feature collapse phenomenon caused by SCL in federated learning, resulting in slow convergence and limited performance improvement.
  • $\bullet$
    We propose a relaxed supervised contrastive loss, which adaptively imposes the divergence penalty on pairs of samples in the same class and prevents the learning of non-distinguishable representations of the samples.
  • $\bullet$
@@
The rest of the paper is organized as follows. We review the prior works in Section 2 and discuss the preliminaries in Section 3. Section 4 presents our approach for federated learning and Section 5 validates its effectiveness. We conclude our paper in Section 6.
@@ 2 Related Works
Federated learning
McMahan et al. [22] proposed FedAvg, a pioneer FL framework that adopts a simple averaging strategy among the locally trained models and uses local SGD steps to alleviate the communication bottleneck. However, it suffers from slow convergence and performance degradation due to the heterogeneous nature of client data in practical use-cases [45]. To address the issue of heterogeneity in FL, numerous approaches have been proposed with two distinct focuses, local training and global aggregation. The major approaches in local training are imposing regularization constraints on model parameters or feature representations. Specifically, they incorporate proximal terms [21], control variates [12, 20], or leverage primal-dual analysis [44, 1] to regularize the weight parameters, while knowledge distillation [18, 40, 15], metric learning [46, 19, 23], logit calibration [42], feature decorrelation [33], or data augmentation [41, 37] are employed for effective representation learning. Our framework also belongs to representation learning, but it particularly focuses on gradient deviations in local training and transferability of trained models across heterogeneous local clients.
Besides the local training methods, server-side optimization techniques have been explored to expedite convergence using momentum [9, 27, 14] or decrease the communication cost by quantization [28, 35, 6, 24]. These server-side works are orthogonal to our client-side approach and integrate well into our framework.
Contrastive learning in FL
Recent works have explored the integration of contrastive learning techniques [25, 3, 7] into federated learning to prevent local client drift and assist local training. FedEMA [46] adopts self-supervised contrastive learning to deal with unlabeled data collected from edge devices. MOON [19] introduces a model-contrastive loss, which aims to align the current local model with the global model while deviating it from the previous local model. FedProc [23] employs a contrastive loss to align local features with the global prototypes to reduce the representation gap, where the global class prototypes are distributed from the server. FedBR [5] conducts contrastive learning to align local and global feature spaces using local and globally shared pseudo-data to reduce local learning bias. In contrast to prior approaches, our framework does not require additional communication overhead for contrastive learning and not rely on global models or prototypes to mitigate the deviations in local training.
@@ 3.1 Problem setup
Suppose that there are $N$ clients, $\{C_{1},...,C_{N}\}=\mathcal{C}$. Each client $C_{i}$ has a dataset $\mathcal{D}_{i}$, which comprises a set of pairs of an example and its class label. The goal of federated learning is to optimize a global model $\theta=[\phi;\psi]$ that minimizes the average losses over all clients as
@@
where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{i}}\left[\ell(\mathbf{x},y;\theta)\right]$ is the empirical loss in $C_{i}$, which is given by the expected loss over all samples in $\mathcal{D}_{i}$, and $\phi$ and $\psi$ denote the parameters of feature extractor and the classifier, respectively. Note that data distributions in individual clients may be heterogeneous, and privacy concerns strictly prohibit the transfer of training data. We employ FedAvg [22] as a baseline algorithm. In the $t^{\text{th}}$ communication round, a central server sends a global model $\theta^{t-1}$ to the active client set $\mathcal{C}_{t}\subseteq\mathcal{C}$. Each client $C_{i}\in\mathcal{C}_{t}$ initializes its parameter $\theta^{t}_{i,0}$ to $\theta^{t-1}$, and performs $K$ iterations for optimization using its local data. The server collects the resulting local models $\theta^{t}_{i,K}$ and computes the global model $\theta^{t}$ for the next round of training by naïvely averaging the local model parameters. This training process is repeated until the global model $\theta^{t}$ converges.
@@ 3.2 Supervised contrastive learning
$\begin{split}\mathcal{L}_{\text{SCL}}(\mathbf{x}_{i},y_{i})=&-{\sum\limits_{y_{j}=y_{i},j\neq i}\left({\langle\phi}(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle/{\tau}\right)}\\ &+\log\Big{(}{\sum\limits_{k\neq i}\exp\left({\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{k})\rangle}/{\tau}\right)}\Big{)},\end{split}$ (5)
which encourages feature representations from the same class to be similar while pushing features from different classes apart.
4 Relaxed Supervised Contrastive Loss
(a) Accuracy
(b) Within-class variance
(c) Between-class variance
(d) Effective rank
(e) Effective rank (centralized setting)
@@ 4 Relaxed Supervised Contrastive Loss
Figure 2: Effects of employing supervised contrastive loss on the CIFAR-100 under non-i.i.d. settings. Black and blue lines denote models trained with $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{SCL}}$, respectively. Dotted and solid lines indicate different data heterogeneity configurations, $\alpha\in\left\{0.05,0.3\right\}$.
@@ 4.1 Benefit of SCL for local training
One of the main challenges in federated learning is inconsistent local updates caused by the heterogeneity of local client data. Zhang et al. [42] present that existing FL methods based on softmax cross-entropy can result in biased local models, and introduce a deviation bound to measure the deviation of the gradient update during the local training. To analyze this further, we revisit the deviation bound in [42] and formulate a sample-wise deviation bound considering all classes as follows.
@@
where $P_{z}^{(y)}=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in\mathcal{O}_{y}}p_{z}(\mathbf{x}_{i})$ is the average prediction score of the samples in a class $y$ to a class $z$, ${\Phi_{y}}=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in O_{y}}{\left\|\phi{(\mathbf{x}_{i})}\right\|_{2}}$ is the average feature norm of the examples in class $y$, and $S_{y}(\mathbf{x})=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in O_{y}}{\langle\phi{(\mathbf{x})},\phi{(\mathbf{x}_{i})}\rangle}$ denotes the average feature similarity to a sample $\mathbf{x}$.
Detailed derivations are provided in the appendix. Eq. (6) presents that the deviation bound depends on the distribution of feature representations, $S_{r}(\mathbf{x})$ and $S_{j}(\mathbf{x})$. When $D(\mathbf{x})\ll 1$, the local updates $\{\Delta\psi_{y}\}_{y\in\mathcal{Y}}$ are prone to deviate from the expected direction, i.e., $\Delta\psi_{r}\phi(\mathbf{x})<0$ and $\Delta\psi_{j}\phi(\mathbf{x})>0$, which incurs inconsistent local training across clients.
We now present how SCL alleviates the deviation of gradient updates.
@@
The optimization of Eq. (4) increases $D(\mathbf{x})$ if $D(\mathbf{x})<\frac{\left(1-{P_{r}^{(r)}}\right){\Phi_{r}|\mathcal{O}_{r}|}}{|\mathcal{Y}|-1}\min\limits_{j\neq r}\left\{\frac{1}{{P}_{r}^{(j)}{{\Phi_{j}}}|\mathcal{O}_{j}|}\right\}$.
Proof. We first derive the lower bound of Eq. (6) as
$D(\mathbf{x})\geq\frac{S_{r}(\mathbf{x})}{\sum\limits_{j\neq r}S_{j}(\mathbf{x})}(1-{P_{r}^{(r)}}){\Phi_{r}|\mathcal{O}_{r}|}\min\limits_{j\neq r}\left\{\frac{1}{{P}_{r}^{(j)}{{\Phi_{j}}}|\mathcal{O}_{j}|}\right\}.$ As the above assumption holds, we formulate the problem as
$\operatorname*{arg\,max}_{\phi}D(\mathbf{x})\approx\operatorname*{arg\,min}_{\phi}\Big{\{}\max\Big{(}0,\sum\limits_{j\neq r}\frac{S_{j}(\mathbf{x})}{|\mathcal{Y}|-1}-S_{r}(\mathbf{x})\Big{)}\Big{\}}.$
By using a smooth approximation to the maximum function with the LogSumExp operator, we derive the upper bound of right-hand side by
@@
This derivation shows the optimization of $\mathcal{L}_{\text{SCL}}$ reduces the upper bound and consequently increases $D(\mathbf{x})$. Refer to the supplementary document for the detailed proof. $\square$
Proposition 1 implies that the SCL objective function mitigates local gradient deviations in the presence of severe data heterogeneity.
4.2 Representation collapse in FL with SCL
Based on our analysis in Section 4.1, we empirically validate the effectiveness of SCL in federated learning under data heterogeneity configurations with Dirichlet parameter $\alpha\in\{0.05,0.3\}$ in Figure 2. We train the ResNet-18 model using the loss function $\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{SCL}}$ at each local client on the CIFAR-100 dataset, using 5% participation rate over 100 distributed clients, where $\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss. As depicted in the figure, while SCL ultimately results in performance improvement by reducing local deviations over the baseline methods, it is accompanied by a noticeable lag in the convergence speed during the early stage of training. We conjecture that, due to limited and skewed local training data, SCL leads to excessively compact representations of the examples in the same classes, hindering effective task transfer across clients in federated learning.
To delve into these phenomena, we first compute the within-class and between-class covariance matrices for the feature embeddings provided by a local model, denoted by $\mathbf{\Sigma}_{W}$ and $\mathbf{\Sigma}_{B}$, respectively. Figure 2 and 2 plot the trace of the two matrices. SCL effectively reduces the within-class variance compared to the baseline model based only on the cross-entropy loss, due to its attraction term between samples from the same class. However, it is noteworthy that SCL also yields a lower between-class variance than the baseline, especially at the early stages of training, despite the repulsion term between the examples in different classes. Since the attraction and repulsion forces interact in contrastive learning, the excessive representation similarity between positive pairs weakens the repulsion force between negative pairs. In other words, the collapse of intra-class representations negatively affects the separation of inter-class examples, leading to an overall reduction in the diversity of the feature representations. To evaluate this feature collapse quantitatively, we observe the effective rank [30] of the covariance matrix of all feature embeddings given by a local model, which measures the effective dimensionality of the learned manifold of training data. Formally, the effective rank is defined as follows:
(a) Between-class variance
(b) Effective rank
(c) Variability collapse index
@@ 4.2 Representation collapse in FL with SCL
Figure 3: Results of our relaxed contrastive learning (RCL) approach, along with other comparisons, on the CIFAR-100 dataset with non-i.i.d. setting ($\alpha=0.05$). RCL improves all metrics over SCL.

### v2

@@ Abstract
We propose a novel contrastive learning framework to effectively address the challenges of data heterogeneity in federated learning. We first analyze the inconsistency of gradient updates across clients during local training and establish its dependence on the distribution of feature representations, leading to the derivation of the supervised contrastive learning (SCL) objective to mitigate local deviations. In addition, we show that a naïve integration of SCL into federated learning incurs representation collapse, resulting in slow convergence and limited performance gains. To address this issue, we introduce a relaxed contrastive learning loss that imposes a divergence penalty on excessively similar sample pairs within each class. This strategy prevents collapsed representations and enhances feature transferability, facilitating collaborative training and leading to significant performance improvements. Our framework outperforms all existing federated learning approaches by significant margins on the standard benchmarks, as demonstrated by extensive experimental results. The source code is available at our project page [^1] https://github.com/skynbe/FedRCL.
@@ 1 Introduction
Federated learning (FL) trains a shared model through the collaboration of distributed clients while safeguarding the privacy of local data by restricting their sharing and transfer. The primary challenge in this learning framework arises from the data heterogeneity across clients and the class imbalance in local data. These problems eventually lead to severe misalignments of the local optima of the client models, hindering the search for better global optima of the aggregated model and slowing down convergence.
To tackle these challenges, most existing approaches focus on minimizing the discrepancy between the global and local models by incorporating regularization techniques on either model parameters [21, 12, 2, 44, 1] or feature representations [19, 23, 18, 40, 15]. However, aligning the local models with the global model entails a trade-off as the global model is not necessarily optimal. Recently, there have been several attempts to analyze the inconsistent local training in a principled way [42, 33]. For example, Zhang et al. [42] investigate the label distribution skewness from a statistical perspective, and introduce a deviation bound for analyzing the inconsistency of gradient updates in local training.
We reformulate the deviation bound proposed in [42] and establish its dependence on the distribution of feature representations. Subsequently, we derive that incorporating the supervised contrastive learning (SCL) objective enhances this bound, resulting in consistent local updates across heterogeneous clients. In other words, we show that employing SCL improves the convergence of federated learning by alleviating the variations of local models.
@@
Figure 1: Performance curves of our framework, dubbed as FedRCL, in comparison to other baselines on the Tiny-ImageNet and CIFAR-100 with non-i.i.d. setting ($\alpha=0.1$). FedSCL incorporates the supervised contrastive learning objective into FedAvg, but it suffers from slow convergence and restrains performance enhancement. Our framework significantly improves both convergence speed and accuracy.
Although SCL is helpful for the optimization in federated learning, the empirical results show that a naïve integration of SCL suffers from slow convergence and limited performance gains as illustrated in Figure 1. Due to the limited and imbalanced training data in a local client, the intra-class attraction force in SCL hampers feature diversity and consequently weakens the transferability of neural networks to diverse tasks. Considering the consolidation principle of federated learning through the aggregation of heterogeneous local models, the lack of transferability impedes the collaborative training process.
To tackle this issue and enhance the transferability of models, we present a novel contrastive learning strategy for federated learning. Our approach imposes the penalty on the sample pairs within the same class that may exhibit excessively high similarity otherwise. Such a simple adaptive repulsion strategy effectively prevents the intra- and inter-class collapse of representations, enhancing the transferability across heterogeneous clients and leading to the discovery of better global optima. Furthermore, we expand the proposed approach to cover all intermediate levels of representations, promoting consistent local updates even further. The proposed approach demonstrates remarkable performance improvements in all datasets and settings consistently, surpassing existing baselines by significant margins. We present the effectiveness and robustness of the proposed method by thorough empirical analysis. Our main contributions are summarized as follows.
  • By reformulating the deviation bound of local gradient update, we theoretically analyze that supervised contrastive learning mitigates inconsistent local updates across heterogeneous clients.
  • We discover the feature collapse phenomenon caused by the standard SCL in federated learning, resulting in slow convergence and limited performance improvement.
  • We propose a relaxed supervised contrastive loss, which adaptively imposes the divergence penalty on pairs of examples in the same class and prevents their representations from being learned to be indistinguishable.
@@
The rest of the paper is organized as follows. We review the prior works in Section 2 and discuss the preliminaries in Section 3. Section 4 presents the proposed approach in the context of federated learning and Section 5 validates its effectiveness empirically. Finally, we conclude our paper in Section 6.
@@ 2 Related Works
This section first overviews the existing FL algorithms, and discusses how contrastive learning has been explored in the context of FL.
2.1 Federated learning
McMahan et al. [22] propose a pioneer FL framework, FedAvg, which aggregates model updates from distributed clients to improve a global model without requiring the exchange of local data. However, it suffers from slow convergence and poor performance due to the heterogeneous nature of client data in practical scenarios [45]. To address the issue of heterogeneity in FL, numerous approaches have been proposed in two distinct directions, local training and global aggregation.
The major approaches in local training are imposing regularization constraints on model parameters or feature representations. Specifically, they incorporate proximal terms [21], introduce control variates [12, 20], or leverage primal-dual analysis [44, 1] to regularize model parameters, while adopting knowledge distillation [18, 40, 15], metric learning [46, 19, 23], logit calibration [42], feature decorrelation [33], or data augmentation [41, 37] for effective representation learning. Our framework also belongs to representation learning, where it particularly focuses on gradient deviations in local training and transferability of trained models across heterogeneous local clients.
Besides the local training methods, server-side optimization techniques have been explored to expedite convergence using momentum [9, 27, 14] or decrease the communication cost by quantization [28, 35, 6, 24]. These server-side works are orthogonal to our client-side approach and are easily combined with the proposed algorithm.
2.2 Contrastive learning in FL
Recent works have explored the integration of contrastive learning techniques [25, 3, 7] into federated learning to prevent local client drift and assist local training. FedEMA [46] adopts self-supervised contrastive learning to deal with unlabeled data collected from edge devices. MOON [19] introduces a model-contrastive loss, which aims to align the current local model with the global model, while pushing the current model away from the local model of the previous round. FedProc [23] employs a contrastive loss to align local features with the global prototypes to reduce the representation gap, where the global class prototypes are distributed from the server. FedBR [5] conducts contrastive learning to align local and global feature spaces using local data and globally shared proxy data to reduce bias in local training. In contrast to prior approaches, our framework does not require additional communication overhead for contrastive learning and does not rely on global models or prototypes to mitigate the deviations in local training.
@@ 3 Preliminaries
Before discussing the proposed approach, we briefly describe the main idea and formulation of federated learning and supervised contrastive learning.
@@ 3.1 Problem setup
Suppose that there are $N$ clients, $\{C_{1},...,C_{N}\}=\mathcal{C}$. Each client $C_{i}$ has a dataset $\mathcal{D}_{i}$, which comprises a set of pairs of an example and its class label. The goal of federated learning is to optimize a global model parametrized by $\theta=[\phi;\psi]$, corresponding to a feature extractor, $\phi$, and a classifier, $\psi$, that minimizes the average losses over all clients as
@@
where $\mathcal{L}_{i}(\theta)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{i}}\left[\ell(\mathbf{x},y;\theta)\right]$ is the empirical loss in $C_{i}$, given by the expected loss over all samples in $\mathcal{D}_{i}$. Note that data distributions in individual clients may be heterogeneous, and privacy concerns strictly prohibit transfering training data across clients. We employ FedAvg [22] as a baseline algorithm. In the $t^{\text{th}}$ communication round, a central server sends a global model $\theta^{t-1}$ to the active client set $\mathcal{C}_{t}\subseteq\mathcal{C}$. Each client $C_{i}\in\mathcal{C}_{t}$ initializes its parameter $\theta^{t}_{i,0}$ to $\theta^{t-1}$, and performs $K$ iterations for optimization using its local data. The server collects the resulting local models $\theta^{t}_{i,K}$ and computes the global model $\theta^{t}$ for the next round of training by simply averaging the local model parameters. This training process is repeated until the global model $\theta^{t}$ converges.
@@ 3.2 Supervised contrastive learning
$\begin{split}\mathcal{L}_{\text{SCL}}(\mathbf{x}_{i},y_{i})=&{\sum\limits_{y_{j}=y_{i},j\neq i}\Big{\{}-\left({\langle\phi}(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle/{\tau}\right)}\\ &+\log\Big{(}{\sum\limits_{k\neq i}\exp\left({\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{k})\rangle}/{\tau}\right)}\Big{)}\Big{\}}.\end{split}$ (3)
This loss function encourages feature representations from the same class to be similar while pushing features from different classes apart.
4 Relaxed Supervised Contrastive Learning
@@
Figure 2: Effects of employing supervised contrastive loss on the CIFAR-100 under non-i.i.d. settings. Black and blue lines denote models trained with $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{SCL}}$, respectively. Dotted and solid lines indicate different data heterogeneity with Dirichlet parameters $\alpha\in\left\{0.05,0.3\right\}$.
@@ 4.1 Benefit of SCL for local training
One of the main challenges in federated learning is inconsistent local updates caused by the heterogeneity of local client data. Zhang et al. [42] present that existing FL methods based on softmax cross-entropy result in biased local models, and introduce a deviation bound to measure the deviation of the gradient update during the local training. To analyze this further, we revisit the deviation bound and formulate a sample-wise deviation bound considering all classes, which is formally defined below.
@@
Let $\mathbf{x}\in\mathcal{O}_{r}$ denote a training example with ground-truth class label $r$. The sample-wise deviation bound is defined as
@@
where $P_{z}^{(y)}=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in\mathcal{O}_{y}}p_{z}(\mathbf{x}_{i})$ means the average prediction score for class $z$, estimated with the examples that belong to class $y$, ${\Phi_{y}}=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in O_{y}}{\left\|\phi{(\mathbf{x}_{i})}\right\|_{2}}$ is the average feature norm of the examples in class $y$, and $S_{y}(\mathbf{x})=\frac{1}{|\mathcal{O}_{y}|}\sum_{i\in O_{y}}{\langle\phi{(\mathbf{x})},\phi{(\mathbf{x}_{i})}\rangle}$ denotes the average feature similarity with respect to an example $\mathbf{x}$.
@@
If $D(\mathbf{x})\ll 1$, the local updates of the parameters in classification layer, $\{\Delta\psi_{y}\}_{y\in\mathcal{Y}}$, are prone to deviate from the desirable direction, i.e., $\Delta\psi_{r}\phi(\mathbf{x})<0$ and $\exists j\neq r$ such that $\Delta\psi_{j}\phi(\mathbf{x})>0$ for $\mathbf{x}\in\mathcal{O}_{r}$.
The proof of Proposition 1 is provided in Section A of the supplementary document. Eq. (4) indicates that the deviation bound of an example depends on the distribution of feature representations with respect to the example, $S_{r}(\mathbf{x})$ and $S_{j}(\mathbf{x})$. This proposition means that lower values of $D(\mathbf{x})$ incur inconsistent local training.
Proposition 1 states that it is possible to prevent the local gradient deviation of each example by increasing $D(\mathbf{x})$. If $\frac{1}{|\mathcal{Y}|-1}\sum\limits_{j\neq r}S_{j}(\mathbf{x})-S_{r}(\mathbf{x})\leq 0$, then the lower bound of $D(\mathbf{x})$ in (4) becomes $\frac{\left(1-{P_{r}^{(r)}}\right){\Phi_{r}|\mathcal{O}_{r}|}}{|\mathcal{Y}|-1}\min\limits_{j\neq r}\left\{\frac{1}{{P}_{r}^{(j)}{{\Phi_{j}}}|\mathcal{O}_{j}|}\right\}$. Thus, we formulate the surrogate objective to minimize $\max\Big{(}0,\frac{1}{|\mathcal{Y}|-1}\sum\limits_{j\neq r}S_{j}(\mathbf{x})-S_{r}(\mathbf{x})\Big{)}$, which is highly correlated to the increase of the lower bound. By using a smooth approximation to the maximum function with the LogSumExp operator, we derive its upper bound as follows
@@
Please refer to Section B for further details. This derivation demonstrates how the optimization of $\mathcal{L}_{\text{SCL}}$ contributes to mitigating local gradient deviations.
@@
Figure 3: Results of our relaxed contrastive learning (RCL) approach on the CIFAR-100 dataset under a non-i.i.d. setting ($\alpha=0.05$). RCL outperform SCL in all metrics.
4.2 Representation collapse in FL with SCL
Based on our analysis in Section 4.1, we empirically validate the effectiveness of SCL in federated learning under data heterogeneity with Dirichlet parameters $\alpha\in\{0.05,0.3\}$ in Figure 2. We train the ResNet-18 model using the loss function $\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{SCL}}$ at each local client on the CIFAR-100 dataset, using 5% participation rate out of 100 distributed clients, where $\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss. As depicted in the figure, while SCL eventually achieves improved performance by reducing local deviations over the baseline methods, it is accompanied by a noticeable lag in the convergence speed during at early stage of training. We conjecture that, due to limited and skewed local training data, SCL leads to excessively compact representations of the examples in the same classes, hindering effective knowledge transfer across clients in federated learning.
To delve into these phenomena, we first compute the within-class and between-class covariance matrices of the feature embeddings provided by a local model, denoted by $\mathbf{\Sigma}_{W}$ and $\mathbf{\Sigma}_{B}$, respectively. Figure 2 and 2 plot the trace of the two matrices. SCL effectively reduces the within-class variance compared to the baseline model only with the cross-entropy loss, due to the attraction term between samples from the same class. However, it is noteworthy that SCL also yields a lower between-class variance than the baseline, especially at the early stages of training, despite the repulsion term between the examples in different classes. Since the attraction and repulsion forces interact in contrastive learning, the excessive representation similarity between positive pairs weakens the repulsion force between negative pairs. In other words, the collapse of intra-class representations negatively affects the separation between inter-class examples, leading to an overall reduction in the diversity of feature representations. To evaluate this feature collapse quantitatively, we observe the effective rank [30] of the covariance matrix of all feature embeddings given by a local model, which estimates the actual dimensionality of the learned feature manifold of training data. Formally, the effective rank is defined as follows: